OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9

OpenStudy (midhun.madhu1987):

Subtract one equation from the other and x will be eliminated. You will be able to find y and then x.

OpenStudy (triciaal):

another option is to solve by substitution from the first equation x = 14 + 2y use this in the second equation (14 + 2y) + 3y = 9 solve for y then x = 14 + 2(-1) = 12